# Remote Sensing Spatio-Temporal Vision-Language Models: A Comprehensive Survey

### v2

@@ V-B3 Dataset Matching Temporal Images, Instruction and Response
This category of datasets aims to align remote sensing temporal images with user instructions and corresponding responses. The instructions are typically expressed in natural language, representing the user’s task requirements. The responses vary depending on the specific task. For the change question answering task, the response is a textual answer. For the visual grounding task, the response may take the form of bounding boxes or segmentation masks. These structured outputs (boxes and masks) can be transformed into serialized text through rule-based methods — a practice that has become increasingly common in the era of multimodal large language models. This unified textual representation facilitates the integration of diverse task formats into a standard VQA-like structure, enabling the joint modeling and execution of multiple tasks within a unified framework.
@@
  • QAG-360K [29]: Designed for Change Question Answering and Grounding, QAG-360K collects bi-temporal image pairs from three change detection datasets including Hi-UCD [231], SECOND [229], and LEVIR-CD [48]. It offers 6,810 image pairs with over 360,000 question–answer–mask triples. Each pair of images contains an average of 53 triples. The images span multiple resolutions (ranging from 0.1m to 3.0m), and cover 24 geographically diverse regions and 10 land-cover categories. Questions are generated by the large language model. Then, a rule-based method leverages annotations of sourced change detection datasets to produce textual answers and corresponding change detection masks. This dataset supports diverse instruction such as identifying the presence, type, magnitude, and ratio of changes.
@@ V-C Comparison of Methods on Benchmark Datasets
To provide a clear and comprehensive overview of existing approaches in the field of spatio-temporal vision-language understanding, we present a comparative analysis of representative methods on two benchmark datasets. These methods span three key categories: change detection, change captioning, and multi-task joint learning of change detection and captioning.
@@ VI Future Prospects and Discussion
Despite the notable progress in remote sensing spatio-temporal vision-language modeling, the field still faces a series of open challenges that hinder the development of truly generalizable, spatio-temporal intelligent systems. This section outlines some promising future research directions that may shape the next stage of this domain.
VI-1 Large-Scale High-Quality Spatio-Temporal Benchmark Datasets
The development of robust spatio-temporal vision-language models critically depends on the availability of large-scale, high-quality datasets. However, existing benchmarks are limited in both spatial coverage and temporal diversity, often focusing on a narrow set of geographic regions or time periods. Furthermore, annotation quality remains a bottleneck: while some studies have attempted to extend dataset size through rule-based or LLM-based automated annotation methods, they frequently suffer from limited textual richness, textual precision, and semantic consistency. Future efforts should emphasize the construction of comprehensive spatio-temporal vision-language datasets that span diverse geographical locations and temporal frequencies (from hours to years). In particular, high-quality natural language annotations aligned with visual spatio-temporal phenomena are needed to support the training and evaluation of more sophisticated models.
VI-2 Multi-modal Temporal Image Vision-Language Understanding
@@
Current spatio-temporal vision-language studies are largely confined to optical imagery, while the remote sensing domain inherently involves multi-modal data sources, including infrared, Hyperspectral, and SAR. These modalities provide complementary views of the same scene under different environmental conditions. Incorporating such data into spatio-temporal vision-language pipelines introduces opportunities for a more robust and comprehensive understanding of dynamic Earth surface changes. However, it also raises fundamental challenges in modality alignment and representation learning. Developing models that can jointly reason over heterogeneous temporal image inputs, while leveraging their unique spectral, geometric, and contextual properties, is an important direction toward building truly generalizable multimodal spatio-temporal systems.
VI-3 Spatio-Temporal Generalization and Variable-Length Reasoning
@@
As the number of satellites with varying revisit cycles increases and temporal image data becomes more abundant, future models must be capable of generalizing across diverse spatio-temporal contexts. This includes handling image sequences of arbitrary temporal length, with varying temporal granularity, resolution, and spatial coverage. Existing approaches often assume fixed-length dual- or tri-temporal image inputs, limiting their applicability to real-world spatio-temporal monitoring tasks. A key research frontier lies in building models that can understand long and variable-length image sequences, capturing both subtle short-term changes and long-term trends. Such capabilities are crucial for spatio-temporal trend analysis, environmental monitoring, and disaster progression modeling, where spatio-temporal dynamics are complex and span multiple scales.
@@
VI-4 Toward Generalizable Spatio-Temporal Vision-Language Reasoning Foundation Models
@@
Despite recent progress in spatio-temporal vision-language models, current approaches remain limited in generalization and reasoning capabilities. Future research should focus on developing unified foundation models capable of generalizing across spatial domains, sensor modalities, temporal scales, resolution levels, and diverse task formats. More importantly, these models should go beyond static perception to establish explicit spatio-temporal reasoning mechanisms—enabling them to infer dynamic patterns, model temporal causality, and perform multi-step reasoning across heterogeneous inputs. Achieving such generalizable and reasoning-capable unified foundation models will be crucial for advancing toward scalable, adaptive, and robust remote sensing systems in real-world, temporally evolving environments.
@@
VI-5 Fine-grained Spatio-Temporal Understanding of Very Large-size Images
@@
Remote sensing imagery inherently captures expansive geographic areas. However, existing models are typically trained and evaluated on relatively small or cropped image patches, restricting their ability to capture both global context and fine-grained local details. Future research should explore efficient and scalable modeling strategies that enable fine-grained understanding over large-scale, multi-temporal image sequences. The key challenge is balancing computational efficiency with perceptual resolution, as directly processing full-resolution sequences is resource-intensive, while downsampling risks losing critical local semantic details. Addressing this challenge is critical for real-world applications such as land use monitoring and urban expansion analysis, where both fine-grained local details and global patterns are essential.
VI-6 Text-driven Temporal Imagery Generation
@@
The scarcity of annotated temporal image sequences remains a major obstacle, particularly for rare events or long-term change phenomena. One promising direction is to explore the text-driven generation of spatio-temporal remote sensing imagery, where large generative models are used to synthesize plausible temporal image sequences conditioned on user instructions or prompts. Such models can serve as virtual data engines, augmenting real-world datasets and enabling hybrid training regimes that combine real and synthetic data. This direction involves maintaining spatial consistency and semantic fidelity over time, which presents unique modeling challenges.
@@ VII Conclusion
This survey presents the first comprehensive review of Remote Sensing Spatio-Temporal Vision-Language Models (RS-STVLMs), a rapidly emerging field that bridges temporal remote sensing imagery and natural language understanding. Moving beyond a task-specific summary, we adopt a systematic perspective: not only outlining the evolution of representative tasks, but also distilling shared key technical components across them. By analyzing key technologies, such as temporal image encoding, vision-language interaction, and language generation, from a unified viewpoint, we uncover generalizable design patterns underlying various RS-STVLM pipelines. This integrative perspective enables a systematic understanding of the RS-STVLMs.
A further distinctive contribution of this work lies in the advanced progress examination of how Large Language Models (LLMs) are reshaping the RS-STVLM research. We review the evolution of LLMs, highlight efficient fine-tuning strategies, and summarize recent advances into three representative LLM-driven paradigms for remote sensing applications.
Moreover, we provide a clear and principled categorization of evaluation metrics and available datasets, tracing their data source, coverage scopes, and benchmarking roles. This structured overview clarifies the fragmented landscape of RS-TVLM benchmarks and offers guidance for future dataset construction and fair model comparison. We also present comparative results across representative methods to illustrate performance trends on benchmark datasets. Finally, we outline key research challenges and promising directions for future research in the RS-STVLM field.
In summary, this survey delivers a holistic, well-organized, and forward-looking review of RS-TVLM research. We strive to chart a clear developmental pathway and highlight key challenges and opportunities, hoping to support and inspire future advances in remote sensing spatio-temporal vision-language understanding.

### v3

@@ V-B3 Dataset Matching Temporal Images, Instruction and Response
This category of datasets aims to align remote sensing temporal images with user instructions and corresponding responses. The instructions are typically expressed in natural language, representing the user’s task requirements. The responses vary depending on the specific task. For the change question answering task, the response is a textual answer. For the visual grounding task, the response may take the form of bounding boxes or segmentation masks. These structured outputs (boxes and masks) can be transformed into serialized text through rule-based methods, a practice that has become increasingly common in the era of multimodal large language models. This unified textual representation facilitates the integration of diverse task formats into a standard VQA-like structure, enabling the joint modeling and execution of multiple tasks within a unified framework.
@@
  • QAG-360K [31]: Designed for Change Question Answering and Grounding, QAG-360K collects bi-temporal image pairs from three change detection datasets including Hi-UCD [287], SECOND [284], and LEVIR-CD [71]. It offers 6,810 image pairs with over 360,000 question-answer-mask triples. Each pair of images contains an average of 53 triples. The images span multiple resolutions (ranging from 0.1m to 3.0m), and cover 24 geographically diverse regions and 10 land-cover categories. Questions are generated by the LLMs. Then, a rule-based method leverages annotations of sourced change detection datasets to produce textual answers and corresponding change detection masks. This dataset supports diverse instruction such as identifying the presence, type, magnitude, and ratio of changes.
@@ V-C Comparison of Methods on Benchmark Datasets
To provide a clear and comprehensive overview of existing approaches in the field of spatiotemporal vision-language understanding, we present a comparative analysis of representative methods on two benchmark datasets. These methods span three key categories: change detection, change captioning, and multi-task joint learning of change detection and captioning.
@@ VI Future Prospects and Discussion
Despite the notable progress in RS-STVLMs, the field still faces a series of open challenges that hinder the development of truly generalizable, spatiotemporal intelligent systems. This section outlines some promising future research directions that may shape the next stage of this domain.
VI-1 Large-Scale High-Quality SpatioTemporal Benchmark Datasets
The development of robust RS-STVLMs relies heavily on the availability of large-scale, high-quality datasets. Existing benchmarks are often limited in geographic regions and temporal diversity. Moreover, annotation remains a bottleneck: while efforts have been made to extend datasets using rule-based or LLM-based automated annotation, challenges persist in ensuring textual richness, precision, and semantic consistency. To address these challenges, future research should focus on constructing comprehensive datasets with a broader geographical range and diverse temporal frequencies, and include high-quality language annotations. One promising approach, inspired by the construction of large-scale static vision-language datasets [204, 203, 311], is to combine large VLMs (e.g., GPT4o) with human-in-the-loop correction. This can leverage the efficiency of AI while ensuring the accuracy and semantic consistency of human annotation.
VI-2 Real-Time Processing and Lightweight Model Design
A major challenge in applying RS-STVLMs to real-world applications is the need for real-time processing of multi-temporal image sequences. To address this, future research should focus on developing lightweight models that retain high accuracy while being computationally efficient. A key strategy for achieving this goal is applying model compression techniques to pre-trained models. Methods such as model pruning [312], quantization [313], and knowledge distillation [314] have proven effective in reducing both model size and inference time, without significantly compromising performance. Besides, exploring structural optimizations in model design is another viable path. For instance, designing sparse attention mechanisms [315] or incorporating efficient state-space models [125] could help reduce model complexity while preserving essential capabilities for spatiotemporal reasoning. By combining these techniques, it is possible to develop models that are not only efficient but also capable of real-time processing in resource-constrained environments.
VI-3 Multi-modal Temporal Image Vision-Language Understanding
@@
Currently, most spatiotemporal vision-language studies focus primarily on optical imagery, while remote sensing inherently involves multi-modal data, including infrared, hyperspectral, and SAR images. These modalities provide complementary information about the same scene under different conditions. Incorporating multi-modal data into RS-STVLMs presents opportunities for more comprehensive understanding of dynamic Earth surface changes. However, integrating such heterogeneous data introduces significant challenges, such as modality alignment and joint representation learning. Future research should focus on cross-modal attention and modality-specific experts [316], which can effectively merge data from different modalities in shared embedding spaces. This would enable the models to leverage the spectral, geometric, and contextual properties of each modality, improving robustness and generalizability.
VI-4 SpatioTemporal Generalization and Variable-Length Reasoning
@@
As the number of satellites with varying revisit cycles increases and temporal image data becomes more abundant, future models must be capable of generalizing across diverse spatiotemporal contexts. This includes handling image sequences of arbitrary temporal length, with varying temporal granularity, resolution, and spatial coverage. Existing approaches often assume fixed-length dual- or tri-temporal image inputs, limiting their applicability to real-world spatiotemporal monitoring tasks. A key research frontier lies in building models that can understand long and variable-length image sequences, capturing both subtle short-term changes and long-term trends. Although some research has been conducted (e.g., TEOChat [94], EarthDial [141], DVLChat [143]), there is still significant room for improvement in current approaches. Learning video-based VLMs [317] may provide inspiration for this research.
@@
VI-5 Toward Generalizable SpatioTemporal Vision-Language Reasoning Foundation Models
@@
Despite recent progress in spatiotemporal vision-language models, current approaches remain limited in generalization and reasoning capabilities. Future research should focus on developing unified foundation models capable of generalizing across spatial domains, sensor modalities, temporal scales, resolution levels, and diverse task formats. More importantly, these models should go beyond static perception to establish explicit spatiotemporal reasoning mechanisms, enabling them to infer dynamic patterns, model temporal causality, and perform multi-step reasoning across heterogeneous inputs. Achieving such generalizable and reasoning-capable unified foundation models will be crucial for advancing toward scalable, adaptive, and robust remote sensing systems in real-world, temporally evolving environments. A promising approach is to build LLM-based foundation models, and use techniques such as reinforcement learning [318] and chain of thought [319] to improve the model’s reasoning ability.
@@
VI-6 Fine-grained SpatioTemporal Understanding of Very Large-size Images
@@
Remote sensing imagery inherently captures expansive geographic areas. However, current models are typically trained and evaluated on smaller or cropped image patches, limiting their ability to capture both global context and fine-grained local details. While there has been some progress in the study of large images in the context of single-image VLMs [253, 320, 321], research on large-size image processing in spatiotemporal VLMs remains limited. Future research should explore efficient and scalable modeling strategies that enable fine-grained understanding over large-scale, multi-temporal image sequences. The key challenge is balancing computational efficiency with perceptual resolution, as directly processing full-resolution sequences is resource-intensive, while downsampling risks losing critical local semantic details. Thus, developing techniques to effectively manage large images while maintaining both spatial and temporal context will be crucial for the next generation of RS-STVLMs.
VI-7 Text-driven Temporal Imagery Generation
@@
The scarcity of annotated temporal image sequences remains a major obstacle, particularly for rare change phenomena. One promising direction is to explore the text-driven generation of temporal remote sensing imagery, where large generative models are used to synthesize plausible temporal image sequences conditioned on user instructions or prompts [92, 93]. Such models can serve as virtual data engines, augmenting real-world datasets and enabling hybrid training regimes that combine real and synthetic data. This direction involves maintaining spatial consistency and semantic fidelity over time, which presents unique modeling challenges. Extending pre-trained generative foundation models such as Text2Earth [204] from single images to spatiotemporal image generation would be a promising approach.
@@ VII Conclusion
This survey presents the first comprehensive review of Remote Sensing SpatioTemporal Vision-Language Models (RS-STVLMs), a rapidly emerging field that bridges temporal remote sensing imagery and natural language understanding. Moving beyond a task-specific summary, we adopt a systematic perspective: not only outlining the evolution of representative tasks, but also distilling shared key technical components across them. By analyzing key technologies, such as temporal image encoding, vision-language interaction, and language generation, from a unified viewpoint, we uncover generalizable design patterns underlying various RS-STVLM pipelines. This integrative perspective enables a systematic understanding of the RS-STVLMs.
A further distinctive contribution of this work lies in the advanced progress examination of how LLMs are reshaping the RS-STVLM research. We review the evolution of LLMs, highlight efficient fine-tuning strategies, and summarize recent advances into three representative LLM-driven paradigms for remote sensing applications.
Moreover, we provide a clear and principled categorization of evaluation metrics and available datasets, tracing their data source, coverage scopes, and benchmarking roles. This structured overview clarifies the fragmented landscape of RS-STVLM benchmarks and offers guidance for future dataset construction and fair model comparison. We also present comparative results across representative methods to illustrate performance trends on benchmark datasets. Finally, we outline key research challenges and promising directions for future research in the RS-STVLM field.
In summary, this survey delivers a holistic, well-organized, and forward-looking review of RS-STVLM research. We strive to chart a clear developmental pathway and highlight key challenges and opportunities, hoping to support and inspire future advances in remote sensing spatiotemporal vision-language understanding.